Pick one publicly traded company and go their website and obtain the most recent annual income statement and balance sheet.  Your paper will explain these two statements in terms of what you have learned so far this semester.  Suggested topics to cover:

Total Revenue, Gross Margin and Net Income


Earnings per share


Total Assets and total equity


Percent of debt to total equity


Current Ratio


Inventory turnover ratio


Receivables turnover ratio


Days Sales Outstanding


Method of Inventory costing and valuation


How does this company’s ratios compare to its peers?


Is this company’s performance improving, staying the same or getting worse?


Also spend some time talking about what this company does…Pick a company that you like their product or service.


Based on your analysis would you invest in this company?


must be 3 pages
Answer #1

We have picked the company Facebook Inc.

The most recent Financial Statement that we get from the website of facebook is for 2018.

All figures are in ($' million)

Total Revenue = 55838

Gross Margin= Total Revenue - Cost of goods sold

we take cost of revenue instead of cost of goods sold as our company is in service sector.

= 55838 - 9355

= 46483

Net Income = 22112

Earnings per share

Basic EPS = 7.65 per share

Diluted EPS = 7.57 per share

Total Assets and total equity

Total Assets = 97334

Total Equity = 84127

Percent of debt to total equity

= Total Liabilities/ Total Equity

=13207/84127

=15.7%

Current Ratio

= Current Asssets / Current liabilities

=50480/7017

=7.19 times

Inventory turnover ratio

= Total revenue/ average inventory

Facebook does not hold any inventory, thus inventory turnover can not be calculated.

Receivables turnover ratio

= Credit revenue/ Average receivables

=55838/6709.50

8.32 times

Note: Assuming total revenue is credit revenue

Average receivable = (Opening receivables + closing receivable)/2

=(5832 + 7587)/2

=6709.50

Days Sales Outstanding

=(Accounts receivable ÷ Annual revenue) × Number of days in the year

=(7587 ÷ 55838) × 365

=49.59 days

Method of Inventory costing and valuation

Facebook does not hold any inventory, thus inventory costing and valuation not required.

How does this company’s ratios compare to its peers?

We can easily compare the ratios of this company to its peer.

The ratios we calculated above are EPS, Current ratio, receivable turnover ratio - these ratios are higher the better.

where as Debt to total equity, Days sales outstanding - these ratios are lower the better

Is this company’s performance improving, staying the same or getting worse?

On comparing with last year data company's performance is improving. There is an increase in Total revenue, Gross margin, Net Profit. however the gross margin ratio has decreased from last year. But the overall performance is improving.

Based on your analysis would you invest in this company?

Companies current stock price is $202.00

companies P/E is 202/7.65

=26.40 times

on the basis of my analysis I would like to invest in the company